# An image speaks a thousand words, but can everyone listen? On image transcreation for cultural relevance

### v2

@@ Abstract
Given the rise of multimedia content, human translators increasingly focus on culturally adapting not only words but also other modalities such as images to convey the same meaning. While several applications stand to benefit from this, machine translation systems remain confined to dealing with language in speech and text. In this work, we take a first step towards translating images to make them culturally relevant. First, we build three pipelines comprising state-of-the-art generative models to do the task. Next, we build a two-part evaluation dataset – (i) concept: comprising 600 images that are cross-culturally coherent, focusing on a single concept per image; and (ii) application: comprising 100 images curated from real-world applications. We conduct a multi-faceted human evaluation of translated images to assess for cultural relevance and meaning preservation. We find that as of today, image-editing models fail at this task, but can be improved by leveraging LLMs and retrievers in the loop. Best pipelines can only translate 5% of images for some countries in the easier concept dataset and no translation is successful for some countries in the application dataset, highlighting the challenging nature of the task. Our code and data is released here. [^1] https://github.com/simran-khanuja/image-transcreation
@@
Note: This paper contains examples of potentially offensive generated images.
@@ 1 Introduction
Since the time ancient texts were first translated, philosophers and linguists have highlighted the need for cultural adaptation in translation processes Jerome (384); Khaldun (1377); Dryden (1694); Jakobson (1959); Nida (1964) – achieving the same “effect” on the target audience is essential Nida (1964). Further, with increased consumption and distribution of multimedia content, scholars in translation studies Chaume (2018); Ramière (2010); Sierra (2008) challenge the notion of simply translating words, highlighting that visuals, music, and other elements contribute equally to meaning. While each modality carries its own information, interaction between modalities creates deeper, emergent meanings. Partial translation disturbs this multimodal interaction and causes cognitive dissonance to the receptor Esser et al. (2016). Traditionally, translation has been associated with language in speech and text. To broaden its scope to all modalities, and emphasize on the translator’s creative role in the process, the term transcreation is seeing widespread adoption today.
@@
Contribution 1 (Task): In this paper, we take a first step towards transcreation with machine learning systems, by assessing capabilities of generative models for the task of image transcreation across cultural boundaries. In text-based systems alone, models struggle with translating culture-specific information, like idioms (Liu et al., 2023). Moreover, to our knowledge, automatically translating visual content has previously been unaddressed.
Contribution 2 (Pipelines): In §2, we introduce three pipelines for this task – a) e2e-instruct (instruction-based image-editing): that edits images directly following a natural language instruction; b) cap-edit (caption -> LLM edit -> image edit): that first captions the image, makes the caption culturally relevant, and edits the original image as per the culturally-modified caption; and c) cap-retrieve (caption -> LLM edit -> image retrieval): that uses the culturally-modified caption from cap-edit to retrieve a natural image instead.
Contribution 3 (Evaluation dataset): Given the unprecedented nature of this task, the evaluation landscape is a blank slate at present. We create an extensive and diverse evaluation dataset consisting of two parts (concept and application), as detailed in §3. Concept comprises 600 images across seven geographically diverse countries: Brazil, India, Japan, Nigeria, Portugal, Turkey, and United States. Five culturally salient concepts and related images are collected across a consistent set of universal categories (food, beverages, celebrations, and so on) making this dataset cross-culturally comparable. Application comprises 100 images curated from real-world applications like educational worksheets and children’s literature.
Figure 2: Pipelines to translate images: e2e-instruct takes as input the original image and a natural language instruction; cap-edit is a modular approach which first captions the image, uses a LLM to edit the caption for cultural relevance, and edits the original image using the LLM-edit as instruction; cap-retrieve similarly captions the image and LLM-edits the caption, but retrieves from a country-specific image dataset instead.
Contribution 4 (Human evaluation): In §4, we conduct human evaluation of images translated for both concept and application, across all seven countries. We find that as of today, image-editing models fail at this task, but can be improved by leveraging LLMs and retrievers in the loop. Even the best models can only successfully translate 5% images for Nigeria in the simpler concept dataset and no image translation is successful for some countries in the harder application dataset. We hope our work paves the way to building culturally inclusive multimodal systems.
@@ 2 Pipelines for Image Transcreation
We introduce three pipelines for image transcreation comprising of state-of-the-art generative models. An overview of each pipeline is in Figure 2.
@@
First, we use out-of-the-box instruction-based image editing models to translate the image in one pass. Specifically, we use InstructPix2Pix Brooks et al. (2023), a model that allows users to define edits using natural language, as opposed to other models requiring text labels, captions, segmentation masks, example output images and so on.
@@
Our second approach is a modular pipeline that offloads some of the requirement of cultural understanding from image editing models to large language models (LLMs). Large language models have been trained on trillions of tokens of text Touvron et al. (2023); Achiam et al. (2023), and have been shown to exhibit at least a certain degree of cultural awareness Arora et al. (2022). Concretely, we adopt a method that first performs image captioning, translates the caption for cultural relevance using an LLM, and then edits the image using a caption-based image editing model. In experiments, we use BLIP2-FlanT5-XXL [^3] https://huggingface.co/Salesforce/blip2-flan-t5-xxl (Li et al., 2023) as the image captioner, GPT-3.5444https://platform.openai.com/docs/models/gpt-3-5 for caption transformation, and PlugnPlay as the image editing model Tumanyan et al. (2023).
@@
2.3 cap-retrieve: Caption, Text-edit, Image Retrieval
In Step-3 of cap-edit, we edit the original image, which can lead to the final output not being reflective of how the concept naturally appears in the target country (§A.2). Hence, here we rely on retrieval from a country-specific image database instead. Concretely, we first caption the image and edit the caption for cultural relevance, similar to cap-edit. Next, we use the LLM-edited caption to query country-specific subsets of LAION Schuhmann et al. (2022). These subsets are created by parsing image URLs and categorizing them based on the country-code top-level domain they contain. For example, URLs featuring “.in” are assigned to the India subset, those with “.jp” are grouped into the Japan subset, etc.
@@ 3 Evaluation Dataset
We design a two-part dataset where the first part (concept) is meant to serve as a research prototype, while the second (application) is grounded in real-world applications like those in Figure 1.
@@ 3.1 Concept dataset
The first part contains images collected for the same universal categories, across seven countries, as shown in Figure 3. We follow the annotation protocol of MaRVL Liu et al. (2021) which involves people local to a region driving the entire annotation process. This approach ensures that the data accurately captures their lived experiences. Concretely, the collection process is as follows:
Country Selection: We select seven geographically diverse countries: Brazil, India, Japan, Nigeria, Portugal, Turkey, and United States. But do geographic borders dictate cultural ones? Cultures constantly change and are hybrid at any point in time Hall (2015). However, audiovisual adaptation is most often equated with national boundaries Moran (2009); Keinonen (2016), given the significant influence of history, policy, and state regulations on media consumption within countries Steemers and D’Arma (2012). Further, from a practical perspective, machine learning systems need data, whose source can be geographically tagged and segregated. While the ultimate goal is to adapt to individual experiences that shape cultural contexts, focusing on the national level serves as a practical starting point, aligning with established practices in related fields.
Category Selection: Ideally, datasets for different cultures should reflect the most salient concepts and their typical visual denotations, while retaining some thematic coherence for comparability. Hence, following Liu et al. (2021), we opt for a list of universal concepts that are cross-culturally comprehensive, as laid out in the Intercontinental Dictionary Series Key and Bernard Comrie (2015).
@@ 3.2 Application dataset
Education: Research suggests that incorporating objects in a child’s surrounding and grounding content in their culture aids learning Council et al. (2015). Looking at math worksheets for grades 1-3, we find this to be true. Hence, we source worksheets from K5 Learning, [^5] https://www.k5learning.com/free-worksheets-for-kids We obtain permission to use and distribute the worksheets for non-commercial research purposes from the publisher. a US-based learning platform. The translation process is tied to the task here, and may not be as straightforward as replacing currency notes in Figure 1. For example, in the left below, the model must find differently-colored elements while retaining the count of each colored object during translation, or on the right, where its necessary to find objects that can be measured using the chosen replacement for a matchstick.
@@
Literature: We curate images from Bloom Library, [^6] https://bloomlibrary.org/ a digital library of stories for children. Specifically, we use the dataset released by Leong et al. (2022) on HuggingFace.  [^7] https://huggingface.co/datasets/sil-ai/bloom-vist A story is sequential, consisting of at least 14-15 images with text. However, dealing with consistency for sequential data is a hard problem in the image-generation space and out-of-scope of the current work. Further, models today struggle to translate a single image alone, as noted in Section 4. Hence, we simply collect the first image in each story along with its text, which is used to guide the translation process. We manually select roughly 60 images out of 400 from the eng subset, making sure the selected images are of high quality and de-duplicated (Figure 5).
@@ 3.3 Why the two-part dataset?
Even though our eventual goal is to translate images for real-world applications, real-world scenes are complex, comprising of multiple interacting objects, and have application-specific constraints, making the task harder. For example, in Figure 4(b), one is constrained to find objects of a specific length that can be measured using a matchstick.
With concept, we build a prototype which has the following features: a) diverse: images are collected across 7 geographically spread-out countries; b) cross-culturally comparable: concepts selected from a universal set of semantic categories (e.g., food, agriculture), hence enabling cross-cultural comparisons; c) single concept or object per image: making it easier to analyse model errors when one image represents a concept in isolation; d) loose constraints on output: the goal is simply to increase cultural relevance while staying within bounds of the universal category.
@@ 4 Human Evaluation and Quantitative Metrics
Evaluation of image-editing models typically relies on quantitative metrics and qualitative analysis of a few select samples. [^8] Some skip a quantitative evaluation altogether as in Hertz et al. (2022). While image-editing focuses on image quality and how closely the edit follows the instruction, image-translation comes with additional requirements such as cultural relevance, meaning preservation, and so on. Hence, we design an extensive questionnaire and conduct human evaluation to assess the quality of all generated images, across both parts of the dataset. Evaluators are shown the source image and the three pipeline outputs in a single instance, (Figure 8). This ensures that scores capture relative differences across pipelines. Further, the order of pipeline outputs is randomized so as to not bias the ratings.
All images are run through all pipelines, separately for each of the 7 chosen countries as target. We ask six questions per instance for concept, and three per instance for application, as detailed below. To put all of this in perspective, 12,500 questions have been given a 5-point rating across 2,400 images, for each country (split across five participants) in the evaluation.
@@
4.1 Questions and Findings: Concept
The translation here is open-ended; our end goal is to make an edit that doesn’t change the universal category, and increases cultural relevance. However, some applications have additional constraints like maintaining the spatial layout (ex: AV media like Doraemon). Hence, our evaluation is multi-dimensional, analysing the strengths and weaknesses of each pipeline, which can help inform its appropriateness for an end-application. We ask human evaluators to give ratings on a 5-point scale on six questions as shown in Table 1. We summarize key findings below, and a detailed analysis for all questions is in §D.
C0: visual-change
First we ask whether the image has been edited at all, to help understand if the edits make sense in the questions that follow. Across all countries, cap-retrieve maximally edits images, with roughly 90% scoring 5 (Figure 6). This is expected since here the original image is not input at all in producing the final image. e2e-instruct on the other hand makes no edit sometimes, with 40-60% images being given a score of 1. For countries like Brazil and US, this pipeline overwhelmingly paints the image with the flag or flag colors (§A.1), explaining the relatively lower number of 1s.
C1: semantic-equivalence
– First, we filter out images with scores 1 or 2 on C0 since here we want to capture that if an edit has been made, is it a meaningful one? In Figure 6, we observe that cap-edit scores highest, while cap-retrieve’s performance varies based on the country. For example, it has higher performance for India and US but low for Nigeria. Image generation models have been shown to be reflective of cultures in the US followed by India for under-specified inputs Basu et al. (2023), which indicates higher quality model representations and training data for these countries, explaining why cap-retrieve would perform better for them.
Quantitative Metrics for image-editing typically capture how closely the edited image matches – (i) the original image; and (ii) the edit instruction. Following suit, we calculate two metrics: a) image-similarity: we embed the original image and each of the generated images using DiNO-ViT Caron et al. (2021) and measure their cosine similarity; and b) country-relevance: we embed the text – This image is culturally relevant to {COUNTRY}, and the edited images using CLIP Radford et al. (2021) and calculate their cosine similarity. We present results for both metrics in Figures 17 and 18. A discussion on correlation of these metrics with human evaluation is in §C.
We find that overall for image-similarity, e2e-instruct scores highest, closely followed by cap-edit, while cap-retrieve lags behind, consistent with human ratings for C0: visual-change and C2: spatial layout. For the country-relevance score, we observe a similar trend as that for C3: cultural-relevance.
Figure 7: Human ratings for the application dataset: Our goal is to test whether the edited image can be used for the application as before (E/S1), and whether it increases cultural relevance (E/S2). We plot the count of images that can do both above (E/S1+E/S2), and observe that even the best pipeline cannot translate any image successfully in some cases, like for Portugal.
@@ 4.2 Questions and Findings: Application
Here, we must make sure that: a) conditioned on the task, the "meaning" of the original image is preserved; b) the cultural relevance increases. Our observations here are as follows:

### v3

@@ Abstract
Given the rise of multimedia content, human translators increasingly focus on culturally adapting not only words but also other modalities such as images to convey the same meaning. While several applications stand to benefit from this, machine translation systems remain confined to dealing with language in speech and text. In this work, we introduce a new task of translating images to make them culturally relevant. First, we build three pipelines comprising state-of-the-art generative models to do the task. Next, we build a two-part evaluation dataset – (i) concept: comprising 600 images that are cross-culturally coherent, focusing on a single concept per image; and (ii) application: comprising 100 images curated from real-world applications. We conduct a multi-faceted human evaluation of translated images to assess for cultural relevance and meaning preservation. We find that as of today, image-editing models fail at this task, but can be improved by leveraging LLMs and retrievers in the loop. Best pipelines can only translate 5% of images for some countries in the easier concept dataset and no translation is successful for some countries in the application dataset, highlighting the challenging nature of the task. Our code and data is released here. [^1] https://github.com/simran-khanuja/image-transcreation
@@ 1 Introduction
Since the time ancient texts were first translated, philosophers and linguists have highlighted the need for cultural adaptation in the process Jerome (384); Khaldun (1377); Dryden (1694); Jakobson (1959); Nida (1964) – achieving the same “effect” on the target audience is essential Nida (1964). Further, with increased consumption and distribution of multimedia content, scholars in translation studies Chaume (2018); Ramière (2010); Sierra (2008) challenge the notion of simply translating words, highlighting that visuals, music, and other elements contribute equally to meaning. While each modality carries its own information, interaction between modalities creates deeper, emergent meanings. Partial translation disturbs this multimodal interaction and causes cognitive dissonance to the receptor Esser et al. (2016). Traditionally, translation has been associated with language in speech and text. To broaden its scope to all modalities, and emphasize on the translator’s creative role in the process, the term transcreation is seeing widespread adoption today.
@@
Contribution 1 (Task): In this paper, we take a first step towards transcreation with machine learning systems, by assessing capabilities of generative models for the task of image transcreation across cultural boundaries. In text-based systems alone, models struggle with translating culture-specific information, like idioms (Liu et al., 2023). Moreover, to our knowledge, automatically transcreating visual content has previously been unaddressed.
Contribution 2 (Pipelines): In §2, we introduce three pipelines for this task – a) e2e-instruct (instruction-based image-editing): that edits images directly following a natural language instruction; b) cap-edit (caption $\rightarrow$ LLM edit $\rightarrow$ image edit): that first captions the image, makes the caption culturally relevant, and edits the original image as per the culturally-modified caption; c) cap-retrieve (caption $\rightarrow$ LLM edit $\rightarrow$ image retrieval): that uses the culturally-modified caption from cap-edit to retrieve a natural image instead. We also experiment with GPT-4o and DALLE-3 to generate new images using culturally-modified captions (§A.4).
Contribution 3 (Evaluation dataset): Given the unprecedented nature of this task, the evaluation landscape is a blank slate at present. We create an extensive and diverse evaluation dataset consisting of two parts (concept and application), as detailed in §3. Concept comprises 600 images across seven geographically diverse countries: Brazil, India, Japan, Nigeria, Portugal, Turkey, and United States. Five culturally salient concepts and related images are collected across a consistent set of universal categories (like food, beverages, celebrations, and so on) from each country. Application comprises 100 images curated from real-world applications like educational worksheets and children’s literature.
Figure 2: Pipelines to transcreate images: e2e-instruct takes as input the original image and a natural language instruction; cap-edit first captions the image, uses a LLM to edit the caption for cultural relevance, and edits the original image using the LLM-edit as instruction; and cap-retrieve uses this LLM-edit to retrieve a natural image from a country-specific image dataset. Given the unprecedented nature of this task, we create pipelines using pre-existing SOTA models, and benchmark them on our newly created test set.
Contribution 4 (Human evaluation): In §4, we conduct human evaluation of images transcreated for both concept and application, across all seven countries. We find that as of today, image-editing models fail at this task, but can be improved by leveraging LLMs and retrievers in the loop. Even the best models can only successfully transcreate 5% images for Nigeria in the simpler concept dataset and no image transcreation is successful for some countries in the harder application dataset.
@@ 2 Pipelines for Image Transcreation
We introduce three pipelines for image transcreation comprising of state-of-the-art generative models. The code to run all pipelines with exact prompts used can be found in Table D.2. An overview of each pipeline is in Figure 2.
@@
First, we use out-of-the-box instruction-based image editing models to translate the image in one pass. Specifically, we use InstructPix2Pix Brooks et al. (2023), a model that allows users to define edits using natural language, as opposed to other models requiring text labels, captions, segmentation masks, example output images and so on. [^3] https://www.timothybrooks.com/instruct-pix2pix
@@
Our second approach is a modular pipeline that offloads some of the requirement of cultural understanding from image editing models to large language models (LLMs). LLMs have been trained on trillions of tokens of text Touvron et al. (2023); Achiam et al. (2023), and exhibit at least a certain degree of cultural awareness Arora et al. (2022). Concretely, we adopt a method that first performs image captioning, edits the caption for cultural relevance using an LLM, and then edits the image using an instruction-based image editing model. In experiments, we use InstructBLIP-FlanT5-XXL [^4] https://huggingface.co/Salesforce/instructblip-flan-t5-xxl (Li et al., 2023) as the image captioner, GPT-3.5555https://platform.openai.com/docs/models/gpt-3-5 for caption transformation, and PlugnPlay as the image editing model Tumanyan et al. (2023).
@@
2.3 cap-retrieve: Caption, edit, retrieve
In cap-edit, the final output is sometimes not reflective of how the concept naturally appears in the target country, due to image-editing models being trained to strictly preserve spatial layout (§A.2). Hence, here we rely on retrieval from a country-specific image database instead. Concretely, we first caption the image and edit the caption for cultural relevance, similar to cap-edit. Next, we use the LLM-edited caption to query country-specific subsets of LAION Schuhmann et al. (2022). These subsets are created by parsing image URLs and categorizing them based on the country-code top-level domain they contain. For example, URLs featuring “.in” are assigned to the India subset, those with “.jp” are grouped into the Japan subset, etc.
@@ 3 Evaluation Dataset
We design a two-part dataset where the first (concept) is meant to serve as a research prototype, while the second (application) is grounded in real-world applications like those in Figure 1.
@@ 3.1 Concept dataset
We collect images for a set of universal categories, across seven countries (Figure 3). We follow the annotation protocol of MaRVL Liu et al. (2021) for which people local to a region drive the entire annotation process, ensuring the collected data accurately captures their lived experiences. Concretely, our collection process is as follows:
Country Selection: We select seven geographically diverse countries: Brazil, India, Japan, Nigeria, Portugal, Turkey, and United States. But do geographic borders dictate cultural ones? Cultures constantly change and are hybrid at any point in time Hall (2015). However, audiovisual adaptation is most often equated with national boundaries Moran (2009); Keinonen (2016), given the significant influence of history, policy, and state regulations on media consumption within countries Steemers and D’Arma (2012). Further, from a practical perspective, ML systems need data, whose source can be geographically tagged and segregated. While the ultimate goal is to adapt to individual experiences that shape cultural contexts, focusing on the national level serves as a practical starting point.
Category Selection: Ideally, datasets for different cultures should reflect most salient concepts as they naturally occur in that culture, while retaining some thematic coherence for comparability Liu et al. (2021). Hence, we opt for a list of universal concepts that are cross-culturally comprehensive, as laid out in the Intercontinental Dictionary Series Key and Bernard Comrie (2015).
@@ 3.2 Application dataset
Education: Research suggests that incorporating objects in a child’s surrounding and grounding content in their culture aids learning Council et al. (2015). Looking at math worksheets for grades 1-3, we find this to be true. We source worksheets from K5 Learning, [^6] https://www.k5learning.com/free-worksheets-for-kids We obtain permission to use and distribute the worksheets for non-commercial research purposes from the publisher. a US-based learning platform. The transcreation process is tied to the task here, and may not be as straightforward as replacing currency notes in Figure 1. For example, in the left below, the model must find differently-colored elements while retaining the count of each colored object during transcreation, or on the right, where its necessary to find objects that can be measured using the chosen replacement for a matchstick.
@@
Literature: We curate images from Bloom Library, [^7] https://bloomlibrary.org/ a digital library of stories for children released for research purposes by Leong et al. (2022). [^8] https://huggingface.co/datasets/sil-ai/bloom-vist Dealing with a sequence of images is out-of-scope of our current work, hence we collect the first image in each story along with its text that is later used to guide the transcreation. We manually select roughly 60 images out of 400 from the eng subset, making sure the selected images are of high quality and de-duplicated (Figure 5).
@@ 3.3 Why the two-part dataset?
Even though our eventual goal is to transcreate images for real-world applications, real-world scenes are complex, comprising of multiple interacting objects, and have application-specific constraints, making the task harder. For example, in Figure 4(b), one is constrained to find objects of a specific length that can be measured using a matchstick.
With concept, we build a prototype which has the following features: a) diverse: images are collected across 7 geographically spread-out countries; b) single concept or object per image: making it easier to analyse model errors when one image represents a concept in isolation; c) loose constraints on output: the goal is simply to increase cultural relevance while staying within bounds of the universal category.
@@ 4 Human Evaluation and Quantitative Metrics
Evaluation of image-editing models typically relies on quantitative metrics and qualitative analysis of a few select samples. [^9] Some skip a quantitative evaluation altogether as in Hertz et al. (2022). While image-editing focuses on image quality and how closely the edit follows the instruction, image-transcreation comes with additional requirements such as cultural relevance, meaning preservation, and so on. Hence, we design an extensive questionnaire and conduct human evaluation to assess the quality of all generated images, across both parts of the dataset (Table 1). Evaluators are shown the source image and the three pipeline outputs in a single instance, (Figure 11). This ensures that scores capture relative differences across pipelines. Further, the order of pipeline outputs is randomized so as to not bias the ratings.
4.1 Questions and Findings: Concept
End Goal: To transcreate the image such that the final image: a) belongs to the same universal category as the original (like food, animals etc.), and b) has higher cultural relevance than the original image, for a given target country.
However, note that we ask many more questions on layout preservation, offensiveness etc, since different applications may have different constraints on the output, as shown in Table 1. A summary of responses are below, while detailed analyses of responses can be found in §D:
C0: Is there any visual change in the generated image, when compared with the source image? cap-retrieve maximally edits images, with roughly 90% scoring 5 (Figure 6); e2e-instruct makes no edit sometimes, with 40-60% images scoring 1; and cap-edit lies mid-way.
C1: If an edit is made, is it meaningful? For images with C0 $>2$, (indicating some visual changes), we observe that cap-edit’s changes maximally retain the universal category, for ex., a food item from country A is changed to another food item from country B; whereas e2e-instruct often makes meaningless edits like pasting flag colors of the target country on the image (§A.1). cap-retrieve is highly variable; for some countries (India, US), it is better than cap-edit and for some (Nigeria), it is very noisy.
C3: Are the edited images more culturally relevant than the original image? Here, we compare the change in the final image’s cultural relevance score with the original image (Figure 6). cap-retrieve has the highest % of images with a positive change, followed by cap-edit after a relatively large gap, while e2e-instruct performs worst. This shows that offloading the cultural translation to LLMs generally helps, and natural images are highly preferred over edited images when assessing for culture.
[…]
@@ 4.2 Questions and Findings: Application
End Goal (Education): To transcreate such that the final image: a) can be used to teach the same concept as the original image (like counting); b) has higher cultural relevance than the original image, for a given target country.
End Goal (Stories): To transcreate such that the final image: a) matches the text of the story; b) has higher cultural relevance than the original image, for a given target country.
Observations: Overall, responses to individual questions are similar to as observed for the concept dataset. The task here is much harder than simply transcreating within a universal category like in concept because of which no image is successfully transcreated by any pipeline for some countries (Portugal). In Figure 7 we see a sample output where e2e-instruct makes the cherries a red that resembles the Japan flag, and cap-edit is a successful transcreation because even though there is a semantic drift from cherries to flowers, the worksheet can be used to teach counting. Detailed results are in §D.1.
(a) original
(b) e2e-instruct
(c) cap-edit
(d) cap-retrieve
Figure 7: Application: Education; Target: Japan — Task: count the number of cherries. cap-edit is a successful transcreation despite the semantic drift from a fruit to a flower, because the final image can be used to teach counting to children.
4.3 Quantitative Metrics
For image-editing, these typically capture how closely the edited image matches – (i) the original image; and (ii) the edit instruction. Following suit, we calculate two metrics:
a) image-similarity: we embed the original image and each of the generated images using DiNO-ViT Caron et al. (2021) and measure cosine similarity
b) country-relevance: we embed the text – This image is culturally relevant to {COUNTRY}, and the edited images using CLIP Radford et al. (2021) and calculate their cosine similarity.
We present results for both metrics in Figures 20 and 21. A discussion on correlation of these metrics with human evaluation is in §C.
We find that overall for image-similarity, e2e-instruct scores highest, closely followed by cap-edit, while cap-retrieve lags behind, consistent with human ratings. However, note that our goal here is to have the right trade-off between image-similarity and the naturalness of the edited image (which cannot be captured by this metric). Figure 8 shows an example of the final image having a high similarity with the original, but nonetheless looks unnatural.
For the country-relevance score, we observe that it has a high recall but low precision. These scores are positively correlated with human ratings for C3: cultural-relevance, but this metric also scores images containing stereotypical artifacts (such as the ones discussed in §A.1) high on cultural relevance.
Our findings above indicate that quantitative metrics cannot sufficiently capture the quality of transcreation of an image, and developing a BLEU-equivalent, but for images, would be necessary to make measurable progress on this task.